OpenStudy (anonymous):

what would be the function and then the description of transformation for g(x)=(x-2)^2

OpenStudy (jdoe0001):

y = A( Bx + C ) + D C = horizontal shift, C>0, to the left, C < 0, to the right

OpenStudy (adilalvi):

polynomial of 2 degree

OpenStudy (anonymous):

so there is no shifting so is there any way you can explain what the graph would look like. Sorry for being so difficult I have multiple graphs to choose from and trying to figure it out

OpenStudy (anonymous):

does it shift up or down man this is confusing to me sorry

OpenStudy (jdoe0001):

parent function is \(\bf g(x)=(x)^2\) then a Shift or Transformation is applied to make it \(\bf g(x)=(x-2)^2\) so it's really a horizontal shift, B = -2, so it's a shift to the right by 2 units of the same parent function :)

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

yw